# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ X-A Preprocessing: Client selection in FL
FedACS [8] is an FA solution to measure the data skewness (severity of data heterogeneity) of the clients. Since data heterogeneity is a significant source of performance degrading of FL tasks, measuring data skewness is proven to be beneficial in assisting client selection for FL. In FedACS, the FL gradients of clients are utilized as FA insights, so that the computation and infrastructure of FL can be reused. The authors prove that the clients with local gradients close to the average gradient tend to have lower data skewness, and are more beneficial to the FL model. The server then formulates a dueling bandit to perform client selection, where clients with lower data skewness are more likely to be selected. In FAVOR [149], the authors formulate a reinforcement learning model running on the server side to perform client selection. In that reinforcement learning model, the states are defined to be a combination of the global model gradient and local gradients of all clients, where the latter is transmitted from the clients and serves as the FA insight. The actions are defined as all possible sets of participating clients in the incoming round. The rewards are defined by the validation accuracy of the model after the training. Such reinforcement learning formulation forces the reinforcement learning model to choose the participating clients that achieve higher validation accuracy. In Oort [150], the authors optimize the training and evaluation of FL models by wisely selecting participating clients. The optimization of training is based on an observation of the authors, that a client inducing higher training loss is more beneficial for the FL training. Therefore, in Oort, the clients upload their training loss (which is a scaler) to the server in each round they participate in, which works as the insight of FA. The server then utilizes a heuristic multi-armed bandit to perform client selection, where clients obtaining higher training losses in previous rounds will be more likely to be selected. In [151], the authors formally prove that selecting clients with higher training loss can accelerate the FL model convergence. Based on the theoretical result, the authors propose a power-of-choice client selection strategy, where a global model is sent to all available clients, the clients upload the training loss as the FA insight, and the server finally selects a set of high-loss clients as the participants in the following several rounds.
X-B Preprocessing: Personalize FL models via client clustering
@@
Among various approaches to personalized FL, clustering-based methods are widely selected for their high reasonability and strong performance. In clustering-based methods, clients are firstly clustered, where those with similar data characteristics are placed in the same cluster. Then, FL is fine-tuned based on the data within one cluster. With such approaches, the personalized models will be trained on sufficient data stored in a number of clients, and tailored for particular data characteristics. The clustering procedure requests the sharing of information about the client data. To preserve privacy, existing clustering-based personalized FL algorithms exactly follow the FA paradigm in clustering the clients. The clustering problems can be regarded as a special kind of clustering problem in general data analytics, and those client clustering algorithms can also be regarded as a special kind of FA-based clustering solutions.
@@
In [9], the authors propose a pioneering clustering personalized FL solution named CFL. In CFL, the cluster assignment of clients is determined by a hierarchical clustering structure: the data distribution similarity of pairs of clients is calculated by the cosine similarity of local gradients; a bipartition of clients is derived by analyzing the similarity matrix; more clusters are derived by hierarchically executing the bipartition algorithm as long as it could improve the overall performance. In [153], the authors propose another pioneering clustering personalized FL solution. It trains a global model with vanilla FL and then uses gradients from all the clients to perform clustering. The clustering is based on the cosine similarity between clients using an agglomerative clustering algorithm. In [154], the authors propose a clustering personalized FL approach by extracting one layer of the gradient uploaded by the clients and clustering clients by calculating the cosine distance of the layer parameters. The authors also propose a decomposition and consolidation scheme to remove some extreme clusters and reassign the clients to other clusters. In [155], the authors propose an Iterative Federated Clustering Algorithm (IFCA) to cluster the clients for personalized FL. In IFCA, the server first formulates one model for each of the clusters and distributes all the models to the participating clients. The clients calculate their local loss of all the models, and one client will be assigned to the cluster where the lowest loss is obtained on the corresponding model. In [156], the authors propose COMET to perform clustered co-distillation in FL. In COMET, there exists a public set of unlabeled data samples. The clients train their personalized model with its local data and then infer the public data to derive the soft decisions. The clients upload their soft decisions on the public data to the server, and the server performs $k$-means on that so that clients with similar soft decisions on the public data will be placed in the same cluster. After that, knowledge sharing is performed in model training. In detail, the clients then train their local model with their local data, while an extra regularization term is applied. The regularization term forces the model to obtain similar soft decisions to other clients in the same cluster on the public data. In [157], the authors consider the problem that the clients may find it hard to cluster into several distinct data distributions. They consider soft clustering, where the distribution of data in each client is considered as a mixture of several clusters. FedSoft is proposed to perform soft clustering by requesting the clients to train with a proximal loss, regulating the training to get close to a mixture of several cluster models. FedSoft system employs an FA subroutine to estimate the weight of each client with all the clustered. Each data sample of one client is evaluated by all the cluster model, and is assigned to the cluster with the smallest loss. The weight of the client is then calculated as the portion of data samples belonging to all the clusters.
X-C Postprocessing: Model evaluation
@@
While the mainstream of FL studies focuses on model training, i.e., deriving a model that is expected to have high performance based on the federated data. Meanwhile, as the evaluation of the trained model also relies on the vast data, leveraging the federated data then becomes an effective approach. In the blog [7] where the term FA is coined, the authors claim that the idea of FA is exactly motivated by the need to evaluate the trained FL model. In Oort [150], the authors propose a system of federated evaluation. Oort automatically adjusts the number of participating clients, so that the accuracy of performance evaluation is guaranteed and the client effort is minimized. Such adjustment is achieved by the confidence bounds derived by Hoeffding’s inequality.
XI Wireless Network Applications
@@
Wireless networks have become a key cornerstone of modern communication systems. Meanwhile, data analytics (machine learning) approaches are widely applied to optimize the performance of wireless network systems. These technologies have been utilized to handle critical tasks such as network slicing, caching management, anomaly detection, and semantic communication. Federated computation, including both FL and FA, is also applied by the researchers to obtain the benefits of privacy preservation, each computing utilization, and communication efficiency. In this section, we focus on the FA solutions applied for the critical optimizations on wireless network systems.
@@
Some FA solutions directly enable non-neural network data analytics powers for wireless network challenges. They design federated versions of data analytics algorithms, such as clustering, to optimize wireless network systems. In [158], an FA solution is proposed for anomaly detection in cellular network antenna tilt. It transforms the data regarding antenna electrical tilt into signature vectors, and then performs federated clustering on these signature vectors, where the local clustering centroids serve as the FA insights. Then, anomaly items are then filtered out by measuring the distance between signature vectors and global clustering centroids. In [159], the problem of energy consumption minimization in mobile edge computing task offloading is considered. To handle the time-varying user task, a federated version of support vector machine (SVM) is proposed. In the federated SVM scheme, the covariance matrix between users is transmitted from server to clients (base stations), and the SVM weights are uploaded from clients to server, and are then aggregated by the server. Eventually, the global SVM can derive the optimal solution of task offloading, with optimized communication and computation overheads. In [160], an FA scheme is utilized in mobile crowdsensing scenarios to discover malicious sensing tasks. In such a scheme, a machine learning module is equipped by each client (detection device). The clients predict whether a task is malicious, and uploads the prediction results to the server as the FA insight. The server aggregates the insights, and derives the updated reputation score of the task proposers to the clients, which helps optimize the prediction results in the future. The global prediction results are also derived by aggregating the client predictions, with extra risk-aware computations.
@@
On the other hand, other FA solutions optimize the wireless network systems based on an FL backbone, i.e., , a neural network is trained collaboratively to directly optimize the system. The FA algorithm then proposes assistance on the FL scheme. Such methodology has been applied in numerous FA research, as is introduced in Section X. In [161], FL is employed to build up a semantic communication system, while FA is applied to assist it via generating the aggregation indicators. The aggregation indicator is generated by the FA local computation scheme, reflecting the suitability between the current channel condition and the local FL model. When the aggregation indicator is received by the server along with the FL model, it is transformed into the weight in weighted FL model aggregation. In [162], the FA-assisted FL scheme is deployed to control the UAV deployment in visible light communications scenarios. It needs to resolve the complex optimization problem jointly considering “UAV deployment, user association, power efficiency, and predictions of the illumination distribution”. To predict the illumination distribution, a convolutional autoencoder is trained by FL methodology. To help the FL training, the FA methodology is also applied. The FA scheme uploads the convolution kernels and bias to the server. By aggregating the convolution kernels, gap matrices can be calculated, which are essential for the evaluation and optimization of the UAV deployment plan. In [163], an FL system is designed to conduct intrusion detection. To tackle the challenge of privacy risk of transmitting gradient, non-IID data, and high communication overhead, the proposed scheme exactly replaces the conventional FL insight derivation and aggregation schemes, by those in FA methodology. Such schemes, utilizing the idea of knowledge distillation, are also termed federated distillation from the FA perspective. In the proposed scheme, the locally generated labels on shared unlabelled data are sent to the server, and then aggregated to form the global hard labels. Then, the server can use the shared data with global labels to train the global model. In [164], an FL model is trained to perform content popularity prediction in Fog-RAN systems. The clustered FL scheme is further deployed based on the federated clustering solution. In the federated clustering scheme, the local features of the clients are computed and uploaded. The server computes the similarity between clients, and splits up new clusters iteratively. Eventually, the final clustering results can be derived, and the personalization of the host FL scheme is improved.
XII Open issues and future directions
XII-A Applications to more complex data scenarios
@@
The fundamental difference between FA and FL lies in the nature of their supporting tasks. The scope of data scenarios studied by existing FA remains small compared with the diversity of data science problems and models in the wild. Designing FA mechanisms to support non-trivial data types, including graph data [165], streaming data[166], key-value data[167], multidimensional data [168], time series data, and important networked applications, including Internet telemetry[169], smart home [170], healthcare [11], web 3.0 and other privacy-critical senarios[171] all present interesting and open challenges.
XII-B Privacy preservation at scale
As FA systems grow in scale with more data or users involved, ensuring privacy becomes more challenging. This challenge is exacerbated when analysts seek high-accuracy analytics. Existing privacy-preserving techniques exhibit distinct characteristics concerning privacy assurance, data utility, and scalability. For example, differential privacy has long been criticized for offering a poor accuracy-privacy tradeoff, prohibiting its applications in domains that request accurate population-level profiling or analytics. Other cryptographic techniques, such as multi-party computation, encounter difficulties in scaling to practical sizes for mobile applications. Consequently, FA mechanisms have to be carefully designed to strike a delicate balance among privacy, utility, and scalability.
XII-C A unified FA framework

### v3

@@ IX-A Preprocessing: Client quality evaluation for FL
Hiding raw data method(s). FedACS [14] is an FA solution to measure the data skewness (severity of data heterogeneity) of the clients. Since data heterogeneity is a significant source of performance degrading of FL tasks, measuring data skewness is proven to be beneficial in assisting client selection for FL. In FedACS, the FL gradients of clients are utilized as FA insights, so that the computation and infrastructure of FL can be reused. The authors prove that the clients with local gradients close to the average gradient tend to have lower data skewness, and are more beneficial to the FL model. The server then formulates a dueling bandit to perform client selection, where clients with lower data skewness are more likely to be selected. In FAVOR [92], the authors formulate a reinforcement learning model running on the server side to perform client selection. In that reinforcement learning model, the states are defined to be a combination of the global model gradient and local gradients of all clients, where the latter is transmitted from the clients and serves as the FA insight. The actions are defined as all possible sets of participating clients in the incoming round. The rewards are defined by the validation accuracy of the model after the training. Such reinforcement learning formulation forces the reinforcement learning model to choose the participating clients that achieve higher validation accuracy. In Oort [87], the authors optimize the training and evaluation of FL models by wisely selecting participating clients. The optimization of training is based on an observation of the authors, that a client inducing higher training loss is more beneficial for the FL training. Therefore, in Oort, the clients upload their training loss (which is a scaler) to the server in each round they participate in, which works as the insight of FA. The server then utilizes a heuristic multi-armed bandit to perform client selection, where clients obtaining higher training losses in previous rounds will be more likely to be selected. In [84], the authors formally prove that selecting clients with higher training loss can accelerate the FL model convergence. Based on the theoretical result, the authors propose a power-of-choice client selection strategy, where a global model is sent to all available clients, the clients upload the training loss as the FA insight, and the server finally selects a set of high-loss clients as the participants in the following several rounds. In CMFL [85], the quality of clients is evaluated as the divergence between the local gradient and global gradient, based on the difference of signs throughout all dimensions of the gradient vector. After that, gradients from those low-quality clients are trimmed. Furthermore, by moving the evaluation procedure from the server to the clients, CMFL reduces the overall communication overhead by preventing the gradient uploads of low-quality clients.
Lessons learned. As the data of participating clients can greatly affect the performance of FL, evaluating the quality of client data becomes a significant applied domain in FA solutions in assisting FL. Existing solutions evaluate client data in many aspects, including the absolute quality and class distribution. These solutions usually recycle intermediate results of FL training to evaluate client data, such as the gradient and the loss. Such an approach has the benefits of computation/communication efficiency, and the inheritance of privacy preservation in FL.
IX-B Preprocessing: Personalize FL models via client clustering
@@
Among various approaches to personalized FL, clustering-based methods are widely selected for their high reasonability and strong performance. In clustering-based methods, clients are firstly clustered, where those with similar data characteristics are placed in the same cluster. Then, FL is fine-tuned based on the data within one cluster. With such approaches, the personalized models will be trained on sufficient data stored in several clients, and tailored for particular data characteristics. The clustering procedure requests the sharing of information about the client data. Similar to those vanilla federated clustering solutions introduced in Section VI-C, the FA schemes for client clustering also just apply the privatization of hiding raw data.
@@
Hiding raw data method(s). In [15], the authors propose a pioneering clustering personalized FL solution named CFL. In CFL, the cluster assignment of clients is determined by a hierarchical clustering structure: the data distribution similarity of pairs of clients is calculated by the cosine similarity of local gradients; a bipartition of clients is derived by analyzing the similarity matrix; more clusters are derived by hierarchically executing the bipartition algorithm as long as it could improve the overall performance. In [89], the authors propose another pioneering clustering personalized FL solution. It trains a global model with vanilla FL and then uses gradients from all the clients to perform clustering. The clustering is based on the cosine similarity between clients using an agglomerative clustering algorithm. In [88], the authors propose a clustering personalized FL approach by extracting one layer of the gradient uploaded by the clients and clustering clients by calculating the cosine distance of the layer parameters. The authors also propose a decomposition and consolidation scheme to remove some extreme clusters and reassign the clients to other clusters. In [90], the authors propose an Iterative Federated Clustering Algorithm (IFCA) to cluster the clients for personalized FL. In IFCA, the server first formulates one model for each of the clusters and distributes all the models to the participating clients. The clients calculate their local loss of all the models, and one client will be assigned to the cluster where the lowest loss is obtained on the corresponding model. In [91], the authors propose COMET to perform clustered co-distillation in FL. In COMET, there exists a public set of unlabeled data samples. The clients train their personalized model with its local data and then infer the public data to derive the soft decisions. The clients upload their soft decisions on the public data to the server, and the server performs $k$-means on that so that clients with similar soft decisions on the public data will be placed in the same cluster. After that, knowledge sharing is performed in model training. In detail, the clients then train their local model with their local data, while an extra regularization term is applied. The regularization term forces the model to obtain soft decisions similar to those of other clients in the same cluster on the public data. In [86], the authors consider the problem that the clients may find it hard to cluster into several distinct data distributions. They consider soft clustering, where the distribution of data in each client is considered as a mixture of several clusters. FedSoft is proposed to perform soft clustering by requesting the clients to train with a proximal loss, regulating the training to get close to a mixture of several cluster models. FedSoft system employs an FA subroutine to estimate the weight of each client with all the clustered. Each data sample of one client is evaluated by all the cluster model and is assigned to the cluster with the smallest loss. The weight of the client is then calculated as the portion of data samples belonging to all the clusters.
Lessons learned. Federated clustering is not only a typical statistical application of FA (discussed in Section VI-C) but also provides valuable optimization on FL. Based on the idea of clustered FL, i.e., clients collaborate with other clients within the same cluster with higher similarity regarding client data. FA solutions for clustering FL derive clusters of clients based on various similarity measures of client data. To ensure privacy preservation, these methods usually reuse the gradient information in FL to represent a fingerprint of client data, so that the privacy level of FL can be preserved.
IX-C Postprocessing: Model evaluation
@@
While the mainstream of FL studies focuses on model training, i.e., deriving a model that is expected to have high performance based on the federated data. Meanwhile, as the evaluation of the trained model also relies on the vast data, leveraging the federated data then becomes an effective approach.
Hiding raw data method(s). In the blog [13] where the term FA is coined, the authors claim that the idea of FA is exactly motivated by the need to evaluate the trained FL model. In Oort [87], the authors propose a system of federated evaluation. Oort automatically adjusts the number of participating clients, so that the accuracy of performance evaluation is guaranteed and the client effort is minimized. Such adjustment is achieved by the confidence bounds derived by Hoeffding’s inequality. In these FA-based model evaluation solutions, the transmitted computation model is usually the learning model, which is consistent with FL. However, the local computation part is conducted by performing model inference, and the derived insight is the performance metric of model inference. It is a simpler form of local computation compared to FL, only requiring the forward propagation phase of FL local computation, and dropping out the more computation-intensive backward propagation phase.
Lessons learned. Evaluating the FL model is the earliest formal application of FA, where the evaluation metrics are computed locally, and aggregated centrally so that the evaluation dataset can be protected. The aggregation of evaluation metrics is usually in the form of mean computation. However, extra issues
X Wireless Network Applications
TABLE XI: Summary of wireless network application.
Reference Type Wireless network application Technical model
Mulvey et al. [96] Wireless network model optimization Anomaly detection Clustering
Wang et al. [93] Wireless network model optimization Energy consumption minimization SVM
Chen et al. [95] Wireless network model optimization Malicious entity detection Customized model
Xing et al. [97] Assisting wireless network FL Semantic communication Customized model
Wang et al. [98] Assisting wireless network FL UAV deployment planning Customized model
Zhao et al. [94] Assisting wireless network FL Intrusion detection Knowledge distillation
Wang et al. [99] Assisting wireless network FL Content popularity detection Clustering
In this section, we focus on the fifth and last class of data analytics tasks in our taxonomy: wireless network applications.
@@ X Wireless Network Applications
Wireless networks have become a key cornerstone of modern communication systems. Meanwhile, data analytics (machine learning) approaches are widely applied to optimize the performance of wireless network systems. These technologies have been utilized to handle critical tasks such as network slicing, caching management, anomaly detection, and semantic communication. Federated computation, including both FL and FA, is also applied by researchers to obtain the benefits of privacy preservation, edge computing utilization, and communication efficiency. In this section, we focus on the FA solutions applied for the critical optimizations on wireless network systems.
Since the “wireless network application” is a relatively large domain, the specific data analytics tasks in FA solutions are diverse and request further classification. In this section, based on the essence of the host data analytics tasks, we further classify the wireless network applications into two types: wireless network model optimization, and assisting wireless network FL. The FA application in wireless networks is relatively primitive, and all existing FA solutions we surveyed do not apply advanced privatization methodology. Instead, they simply apply the methodology of hiding raw data via the federated computation paradigm.
X-A Wireless network model optimization
The complexity of wireless network forces their designers to include many model-based structures to give data-driven intelligence to the wireless network. Some of these models are not as complex as neural networks. Therefore, FA is applied in these tasks as their federated computation solutions to utilize privacy-preserving data. These model-based tasks include SVM [93], clustering [96, 99], and knowledge distillation [94].
@@ X-A Wireless network model optimization
Hiding raw data method(s). In [96], an FA solution is proposed for anomaly detection in cellular network antenna tilt. It transforms the data regarding antenna electrical tilt into signature vectors and then performs federated clustering on these signature vectors, where the local clustering centroids serve as the FA insights. Then, anomaly items are then filtered out by measuring the distance between signature vectors and global clustering centroids. In [93], the problem of energy consumption minimization in mobile edge computing task offloading is considered. To handle the time-varying user task, a federated version of support vector machine (SVM) is proposed. In the federated SVM scheme, the covariance matrix between users is transmitted from the server to clients (base stations), and the SVM weights are uploaded from clients to the server and are then aggregated by the server. Eventually, the global SVM can derive the optimal solution of task offloading, with optimized communication and computation overheads. In [95], an FA scheme is utilized in mobile crowdsensing scenarios to discover malicious sensing tasks. In such a scheme, a machine learning module is equipped by each client (detection device). The clients predict whether a task is malicious, and uploads the prediction results to the server as the FA insight. The server aggregates the insights and derives the updated reputation score of the task proposers to the clients, which helps optimize the prediction results in the future. The global prediction results are also derived by aggregating the client predictions, with extra risk-aware computations.
X-B Assisting wireless network FL
While many wireless network applications utilize non-complex models as introduced in the previous part, others apply neural networks to handle sophisticated tasks faced by wireless networks, and also FL when federated computation paradigm. Similar to those in Section IX, FA solutions are also applied to assist these wireless network FL models, which is another type of FA in wireless network applications
@@ X-B Assisting wireless network FL
Hiding raw data method(s). In [97], FL is employed to build up a semantic communication system, while FA is applied to assist it via generating the aggregation indicators. The aggregation indicator is generated by the FA local computation scheme, reflecting the suitability between the current channel condition and the local FL model. When the aggregation indicator is received by the server along with the FL model, it is transformed into the weight in weighted FL model aggregation. In [98], the FA-assisted FL scheme is deployed to control the UAV deployment in visible light communications scenarios. It needs to resolve the complex optimization problem jointly considering “UAV deployment, user association, power efficiency, and predictions of the illumination distribution”. To predict the illumination distribution, a convolutional autoencoder is trained by FL methodology. The FA methodology is also applied to help with FL training. The FA scheme uploads the convolution kernels and bias to the server. By aggregating the convolution kernels, gap matrices can be calculated, which are essential for the evaluation and optimization of the UAV deployment plan. In [94], an FL system is designed to conduct intrusion detection. To tackle the challenge of privacy risk of transmitting gradient, non-IID data, and high communication overhead, the proposed scheme exactly replaces the conventional FL insight derivation and aggregation schemes, by those in FA methodology. Such schemes, utilizing the idea of knowledge distillation, are also termed federated distillation from the FA perspective. In the proposed scheme, the locally generated labels on shared unlabelled data are sent to the server and then aggregated to form the global hard labels. Then, the server can use the shared data with global labels to train the global model. In [99], an FL model is trained to perform content popularity prediction in Fog-RAN systems. The clustered FL scheme is further deployed based on the federated clustering solution. In the federated clustering scheme, the local features of the clients are computed and uploaded. The server computes the similarity between clients and splits up new clusters iteratively. Eventually, the final clustering results can be derived, and the personalization of the host FL scheme is improved.
We summarize the FA applications in wireless networks in Table XI. Wireless network application refers to their task from the perspective of wireless networks. Technical model refers to the technical/mathematical task from the perspective of data analytics and artificial intelligence. It is marked “customized model” if its model formulation cannot be summarized into any well-known model.
Lessons learned. FA applications on wireless network systems generally follow two types: optimizing some non-neural network models that request private data for functionality or assisting the FL models deployed in wireless network systems. FA helps wireless network systems in many tasks including anomaly detection and all kinds of planning and decision-making. The mathematical formulation of FA solutions may follow some classical problems in data analytics (e.g., SVM and clustering) or include customized analytics tasks that fit the tailored need in wireless network systems.
XI Open Issues and Future Directions
In this section, we discuss some open issues and potential directions for future FA research, so that a bigger picture of future research will be presented. We present seven open issues that cover three directions to improve the research status of FA. In the first direction, FA extends its realm of applications, so that more data scenarios can be supported (Section XI-A), and eventually a unified FA framework can be realized to tackle all kinds of data scenarios (Section XI-B). In the second direction, the privacy preservation of FA could be improved. Privacy preservation can be achieved in large-scale systems (Section XI-C), and different privacy criteria utilized in FA can be measured and compared (Section XI-D). In the last direction, FA is desired to be more convenient for real-world deployment. FA systems are excepted to be proposed with considerations on many systematic factors (Section XI-E); novel technologies in the field of wireless communication are expected to optimize current FA workflow (Section XI-F); all kinds of resources in FA systems are expected to be properly managed (Section XI-G). These open issues and future directions are promising to improve the value of FA to digest valuable information from big data analytics with the strongest respect to user privacy.
XI-A Applications to more complex data scenarios
@@
The fundamental difference between FA and FL lies in the nature of their supporting tasks. The scope of data scenarios studied by existing FA remains small compared with the diversity of data science problems and models in the wild. Designing FA mechanisms to support non-trivial data types, including graph data [192], streaming data[193], key-value data[194], multidimensional data [195], time series data, and important networked applications, including Internet telemetry[196], smart home [197], healthcare [16], web 3.0 and other privacy-critical senarios[198] all present interesting and open challenges. For example, current FA research in graph data focuses on deriving graph metrics from client data with private preservation. However, many real-world graph data are associated with attached data that is not in the form of graphs (e.g., semantic web). How to perform fusion FA that performs analytics with fusion of structural and non-structural data, is a promising and challenging research direction.
XI-B A unified FA framework